OpenStudy (anonymous):

A random sample of 150 students is chosen from a population of 5,000 students. If the mean IQ in the sample is 110 with a standard deviation of 10, what is the 95% confidence interval for the students' mean IQ score?

OpenStudy (amistre64):

how much of this do you already know?

OpenStudy (anonymous):

Well i know most of it kind of, im just not sure how to find the 95% interval

OpenStudy (anonymous):

would it b 104

OpenStudy (anonymous):

No I don't believe so

OpenStudy (amistre64):

\[\Large \frac{x}{n}\pm Z_{.95}\sqrt{\frac{x(n-x)}{n^3}}\]

OpenStudy (anonymous):

I ended up getting a margin of error of 6.7 which doesnt add up to any of the given answers >_<

OpenStudy (amistre64):

or\[\Large \hat p\pm Z_{.95}\sqrt{\frac{\hat p(1-\hat p)}{n}}\]

OpenStudy (amistre64):

im getting a little mixed up between proportion and means :)

OpenStudy (anonymous):

Ugh this is so confusing

OpenStudy (amistre64):

A random sample of 150 students n = 150 is chosen from a population of 5,000 students. irrelevant If the mean IQ in the sample is 110 x bar = 110 with a standard deviation of 10 s = 10 , what is the 95% confidence interval for the students' mean IQ score? z = 1.96; but we might need to use a t score instead?

OpenStudy (amistre64):

\[110\pm1.96\sqrt{\frac{10^2}{150}}\]or change the 1.96 to its t value

OpenStudy (anonymous):

I haven't learning anything about t scores yet o.o

OpenStudy (amistre64):

then its most likely fine as the z score :)

OpenStudy (amistre64):

do you hav a ti83 by chance?

OpenStudy (anonymous):

ti83?

OpenStudy (amistre64):

its a nice calculator is all

OpenStudy (anonymous):

Oh no i wish lol, I only have the calculator on my phone >_<

OpenStudy (amistre64):

our E should be:\[E=1.96\sqrt{\frac{10^2}{150}}\] \[E=1.96\sqrt{\frac{100}{150}}\] \[E=1.96\sqrt{\frac{2}{3}}\] \[E=1.96\sqrt{.66666}\]

OpenStudy (anonymous):

I got 1.6?

OpenStudy (amistre64):

thats good

OpenStudy (amistre64):

so, the interval is 110 +- E

OpenStudy (anonymous):

soo it should be 108.4-111.6?

OpenStudy (amistre64):

well, apart from interval notation ... yes 108.4 to 111.6 is the interval

OpenStudy (anonymous):

okay , thank you for your help!

OpenStudy (amistre64):

youre welcome :)